Re: TOF last night...

Date view Thread view Subject view Author view Attachment view

From: Tonko A. Ljubicic (tonko_at_bnl.gov)
Date: Mon Mar 17 2003 - 09:16:41 EST


On Mon, 17 Mar 2003, W.J. Llope wrote:

>
> ...ran fine. data looks A.O.K.....
>
> looks like eleanor's reseating of a lemo cable at the TOF TCD
> did the trick... thanks again eleanor..

Hm, according to the shift-leader, it did NOT run fine and as such is
NOT in the run... I don't know anything more...

        -- Tonko

--
Tonko (Ante) Ljubicic                         Email: tonko_at_bnl.gov
Bldg. 510A                                    Tel: +1-631-344-7346
Brookhaven National Lab., P.O. Box 5000       Fax: +1-631-344-4206
Upton, NY, 11973

Date view Thread view Subject view Author view Attachment view

This archive was generated by hypermail 2.1.4 : Thu Jul 24 2003 - 00:39:38 EDT